# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v1

@@ 5.4 Effectiveness of C-SAA
This section first compares the performance of the C-SAA method against the three baseline methods. Throughout the following sections, only normalized values of the metrics are reported due to the confidentiality of the data.
5.4.1 Overall Results
Table 3 compares the realized objective value of Problem (18) for C-SAA and the baseline methods. For C-SAA, only the metrics of the best prediction model are recorded. It can be observed that C-SAA consistently has the lowest objective value across different late delivery penalties. C-SAA maintains a lower objective value on average across all $\gamma^{+}$ levels (approximately 0.80), compared to Point-Pred.-Opt. (0.87) and Empirical-SAA (0.83), highlighting its stability and reliability.
A perfect balance between the fulfillment costs and delivery timeliness control needs to be tailored for the specific business needs. On the one hand, a lower fulfillment cost brings mores immediate profits. On the other hand, lower delivery timeliness can significantly increase customer satisfaction, thus retaining more loyal customers, benefiting the company in long-term operations. In CSOFP, the trade-off between these short-term and long-term profits is controlled by the key parameter $\gamma^{+}$, which determines the penalty of late deliveries in terms of costs. The trend indicates that C-SAA’s objective value decreases more sharply compared to other methods as $\gamma^{+}$ increases, making it particularly effective in scenarios where delivery timeliness is crucial. C-SAA outperforms Point-Pred.-Opt. by 2% (0.92 vs. 0.94) when $\gamma^{+}=2.5$ and by 14% (0.63 vs. 0.79) when $\gamma^{+}=40$, indicating that C-SAA’s advantage becomes more pronounced as the penalty increases. While the difference between C-SAA and Empirical SAA in terms of objective value is not huge, the subsequent section will further illustrate the performance gaps between these two methods, as Empirical SAA is more sensitive and could potentially perform worse with larger sample sizes.
Table 3: Simulated Realized Objective Values of C-SAA vs. the Three Baselines
Method $\gamma^{+}=2.5$ $\gamma^{+}=5$ $\gamma^{+}=10$ $\gamma^{+}=20$ $\gamma^{+}=40$
Greedy 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-)
Point-Pred.-Opt. 0.94 $(2.6\times 10^{-4})$ 0.92 $(3.2\times 10^{-3})$ 0.88 $(4.2\times 10^{-3})$ 0.83 $(5.5\times 10^{-3})$ 0.79 $(7.0\times 10^{-3})$
Empirical-SAA 0.93 $(0.9\times 10^{-4})$ 0.91 $(2.0\times 10^{-3})$ 0.86 $(3.1\times 10^{-3})$ 0.76 $(1.0\times 10^{-2})$ 0.64 $(1.2\times 10^{-2})$
C-SAA 0.92 $(1.3\times 10^{-4})$ 0.89 $(3.3\times 10^{-3})$ 0.82 $(5.5\times 10^{-3})$ 0.72 $(1.1\times 10^{-2})$ 0.63 $(1.3\times 10^{-2})$
  • Note: Values represent the normalized mean (standard error) across five different random seeds for the realized deviations simulation. The normalization is performed by setting the value for the Greedy method to 1.
@@ 5.4.2 Value of Contextual Information
To facilitate a closer examination of the performance of these SAA-based models, Figure 4 shows the impact of varying sample sizes. On the one hand, larger sample sizes should provide a more accurate approximation of the expected value in Problem (18), assuming the delivery deviation distributions are well-estimated. On the other hand, it is important to maintain a reasonably small sample size to ensure computational efficiency. For both C-SAA methods, the effect of increasing sample sizes is obvious. The objective values are reduced from 0.86 to 0.82 for C-SAA-CLF and 0.87 to 0.84 for C-SAA-QR on average, indicating their increasing alignment with the true risk-neutral objective function. This improvement is particularly evident in their enhanced effectiveness at avoiding cumulative lateness, defined to be the total number of late delivery days across all units in the orders. The cumulative lateness decreases by 21% for C-SAA-CLF and by 18% for C-SAA-QR.
In contrast, Empirical-SAA struggles to improve its performance in both the realized objective value and the cumulative lateness as the sample sizes grows beyond 10. This can be attributed to the fact that purely using the empirical distribution is not sufficient to capture the varying delivery deviation distributions that may be dependent upon contextual information unique upon each order. A biased distributional forecasts thus guide the model wrongly to create solutions with inferior quality. Based on these observations, both conditional distributional forecasting methods proposed in the paper are successfully in this application, demonstrating a stronger performance compared to using empirical distribution for delivery deviation forecasting. In summary, the value of contextual information is particularly important in order fulfillment decision-making, where the uncertainty of delivery deviation can vary drastically across different states of the environment.
(a) Performance in Realized Objective Value.
@@ 5.4.3 Value of Distributional Forecasts
This section examines the advantage of employing distributional forecasting models over the simpler and more common point prediction model for CSOFP. As supported by theory, simply using the conditional mean estimate of deviations to replace the expected value in CSOFP can introduce bias when approximating the true objective value. This approach may yield worse solutions compared to employing conditional distributional estimators. These performance differences become even more pronounced as late penalties increase, as illustrated in Figure 5. With higher late penalties, accurate forecasting of deviations becomes more crucial; the results reveal the inherent bias of Point-Pred.-Opt, leading to lower quality solutions. This is not only evident in the realized objective value (Figure 5(a)) but also in the cumulative lateness (Figure 5(b)). For example, the cumulative lateness for Point-Pred.-Opt. only improves by 7% from $\gamma^{+}=2.5$ to $\gamma^{+}=40$, whereas C-SAA-CLF and C-SAA-QR show significantly greater improvements of approximately 35% and 32%, respectively. This demonstrates that the C-SAA methods not only maintain lower cumulative lateness overall but also improve more significantly as the penalty for late deliveries increases, further highlighting the benefits of distributional forecasting.
(a) Performance in Realized Objective Value.
@@
Figure 5: C-SAA vs. Point-Pred.-Opt. Across Different Late Delivery Penalties.
5.5 Evaluation of C-RO
This section evaluates the performance of the C-RO methods, whose results are shown in Figure 6. The C-RO methods generally provide more conservative solutions than C-SAA, with C-RO-B showing a performance closer to C-SAA, as shown by the realized objective values in Figure 6(a). Notably, C-RO-B outperforms C-RO-D by offering less conservative yet still robust solutions, while not requiring full distributional information. In real-time fulfillment scenarios, where sampling can be time-prohibitive, especially during peak hours, C-RO-B offers a quick and effective alternative, delivering reasonably high-quality solutions. Additionally, the effectiveness of C-RO methods in controlling late deliveries remains consistent as the penalty increases (see Figure 6(b)). This stability occurs because C-RO methods already account for worst-case scenarios, making them less sensitive to changes in the late delivery penalty, provided that the samples or uncertainty sets encompass these scenarios. As a result, C-RO methods outperform C-SAA in reducing cumulative lateness when the late delivery penalty is low but are outperformed by C-SAA when $\gamma^{+}>10$. However, the C-RO methods incur higher fulfillment costs—calculated as the total shipping costs and fixed non-parcel costs, minus item consolidation discounts across all orders—by up to 20% on average (see Figure 6(c)) This increase is driven by their risk-averse approach, which prioritizes avoiding extreme deviations by opting for more expensive fulfillment options.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
(c) Performance in Fulfillment costs.
Figure 6: Performance of C-RO Methods.
@@ 5.5.1 Flexibility of C-RO-B
The effectiveness of the C-RO models in controlling delivery deviations is evident since they can hedge against the worst-case deviation and potential misspecification of the prediction models. However, their robustness to extreme delivery deviations comes at the cost of more conservative solutions, which can result in higher fulfillment expenses. Fortunately, the C-RO-B models, beyond their distribution-free property, are equipped with the flexibility to control their level of conservativeness. Specifically, they are determined via the coverage of the prediction intervals and the robust uncertainty budget. Smaller prediction interval coverage and uncertainty budget may lead to less conservative solutions, making the models more beneficial in terms of addressing forecasting misspecifications rather than purely managing risk. Conversely, larger coverage and budget aim at minimizing the impact of worst deviations as much as possible with the sacrifice of overall solution quality.
The impact of different prediction interval coverages and robust uncertainty budgets on the quality of solutions obtained from C-RO-B is examined in Figures 7 and 8, respectively. As the prediction interval coverage increases from 50% to 95% in Figure 7(a) , both C-RO-B-CLF and C-RO-B-QR show a noticeable decline in the realized objective value, with the lowest values around 0.85 for C-RO-B-CLF and 0.86 for C-RO-B-QR at 95% and 90% coverage, respectively. This represents a reduction of approximately 6.5% for both C-RO-B methods compared to the objective values at 50% coverage. This suggests that increasing the coverage interval initially benefits the overall performance, up to a point. This is consistent with the general decline trend in cumulative lateness as prediction interval increases (see Figure 7(b)). However, at 99% coverage, there is a sharp increase in the objective value, This trend highlights that while conservative prediction intervals can reduce lateness, they also introduce higher fulfillment costs due to more conservative decisions, thus increasing the overall objective value. For C-RO-B-CLF, this increase in objective value is also due to a declined prediction performance with the largest prediction interval, as indicated by the final increase of cumulative lateness at 99% coverage.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
@@
A similar pattern is observed for varying robust budgets. As $\eta$ increases, the robust budget becomes larger, leading to more conservative solutions. For both C-RO-B methods, larger budget results in a more significant decrease in cumulative lateness (see Figure 8(b)), but the results are less clear when it comes to the objective value (see Figure 8(a)). The realized objective value decreases sharply as the robust budget scaling factor $\eta$ increases from 0.0 to 0.01 for both C-RO-B-CLF and C-RO-B-QR. However, after $\eta=0.01$, there is a noticeable jump in the objective value for C-RO-B-QR, while C-RO-B-CLF continues to stabilize. This jump suggests that C-RO-B-QR adopts more conservative decisions with higher robust budgets, which leads to an increase in fulfillment costs. Therefore, C-RO-B-QR hits its lowest objective value at $\eta=0.01$ whereas C-RO-B-CLF is able to obtain its minimum objective value at $\eta=1$. The analysis reveals a trade-off between increasing robustness and striking a balance between lower fulfillment costs and mitigating extreme deviation risks.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
@@ 5.6 Impact of Consolidation Factor
This section provides a discussion on the impact of item consolidation. CSOFP relies on the coefficients $\beta_{k}$ to control the incentive to batch items. The industrial partner does not know the exact percentage of discount to be offered for item consolidation, which may depend on a variety of convoluted factors. Therefore, this section performs sensitivity analysis on a range of possible discount factors for the CSO models. The consolidation level, calculated as the percentage of units that are batched (i.e., sourced and shipped by the same location-carrier pair) in all orders, is introduced to measure the effectiveness of the methods in identifying item batching opportunities for cost discounts within each order.
@@
Figure 9 illustrates the performance of C-SAA, C-RO-B, and C-RO-D under a series of different discount factors. As shown in Figure 9(a), C-SAA still consistently exhibits the lowest objective value across all discount factors, with C-RO-B comes next. As expected, all three methods exhibit lower objective values due to lower fulfillment costs as the discount factor increases (see Figure 9(b)). This aligns with the intuition that higher discounts encourage more cost-efficient fulfillment strategies.
The decrease in both C-RO methods can be explained by their higher consolidation levels as the discount factors increases. C-RO methods have higher fulfillment costs, so they benefit more from discounts and capitalize on this opportunity.
The trend in consolidation level for C-SAA is less straightforward. While it peaks at a discount factor of 0.25, the consolidation level remains lower than the baseline throughout. This suggests that, despite some consolidation opportunities, the C-SAA model does not exploit them to the same extent as the C-RO models. One possible explanation is that the benefits of consolidation in C-SAA might not be sufficient to outweigh the penalties associated with deviations. The peak at a discount factor of 0.25 could be an anomaly or indicative of specific conditions where consolidation was briefly more favorable, but overall, the model prioritizes minimizing deviation penalties over achieving higher consolidation.
(a) Performance in Realized Objective Value.
(b) Performance in Fulfillment Costs.
(c) Performance in Consolidation Level.
@@
Figure 9: Performance of CSO Models Across Different Discount Factors.

### v2

@@ 6.6 Effectiveness of C-SAA
This section further compares the performance of the C-SAA methods with the two Empirical-SAA baselines and the PTO.
@@ 6.6.1 Value of Contextual Information
To facilitate a closer examination of the performance of the SAA-based models, Figure 4 shows the impact of varying sample sizes ($N_{1}$). On the one hand, larger sample sizes should provide a more accurate approximation of the expected value in Problem (6), assuming the delivery deviation distributions were well-estimated. On the other hand, it is important to maintain a reasonably small sample size to ensure computational efficiency. For both C-SAA methods, the effect of increasing sample sizes is obvious. The average objectives are reduced from 0.85 to 0.82 for C-SAA-CLF and 0.88 to 0.84 for C-SAA-QRF, indicating their increasing alignment with the true risk-neutral objective function. This improvement is particularly evident in their enhanced effectiveness at avoiding cumulative lateness, defined to be the total number of late delivery days across all units in the orders. The cumulative lateness decreases by 14% for C-SAA-CLF and by 12% for C-SAA-QRF.
In contrast, both Empirical-SAA and C-Empirical-SAA struggles to improve there performance in both the average realized objective value and the cumulative lateness as the sample size grows. This is because relying solely on the empirical distribution fails to capture the variability in delivery deviation distributions, which may depend on contextual information specific to each order. A biased distributional forecasts thus guide the model wrongly to create solutions with inferior quality. Based on these observations, both contextual distributional forecasting methods proposed in the paper are successful in this application, demonstrating a stronger performance compared to using empirical distribution for delivery deviation forecasting. In summary, the value of contextual information is particularly important in order fulfillment decision-making, where the uncertainty of delivery deviation can vary drastically across different states of the environment.
(a) Performance in Average Realized Objective Value.
@@ 6.6.2 Value of Distributional Forecasts
This section examines the advantage of employing distributional forecasting models over the simpler and more common point prediction model within the CSO framework. As supported by theory, simply using the conditional mean estimate of deviations to replace the expected value in CSOFP can introduce bias when approximating the true objective value, thus degrading solution quality. Figure 5 confirms that, under changing $\gamma^{+}$, CSO methods with distributional estimator (C‑SAA‑CLF and C‑SAA‑QRF) consistently outperform PTO in both average realized objective value (Figure 5(a)) and cumulative lateness (Figure 5(b)). In fact, the performance differences become more pronounced as late delivery penalty ($\gamma^{+}$) increases.
When $\gamma^{+}$ raises from 2.5 to 40, PTO achieves only a 7% reduction in cumulative lateness, whereas C-SAA-CLF and C-SAA-QRF realize reductions of approximately 22% and 20%, respectively. These results demonstrate that distributional forecasting not only enables lower overall lateness but also leverages higher penalties more effectively, yielding a sharper decline in average objective and making C‑SAA particularly well suited to settings where delivery timeliness is critical.
(a) Performance in Average Realized Objective Value.
@@
Figure 5: C-SAA vs. PTO Across Different Late Delivery Penalties.
6.7 Cost of Robustness
To quantify the trade‑off between robustness and efficiency, C‑RO‑B‑CLF and C‑RO‑D‑CLF were compared with C‑SAA‑CLF as the late‑delivery penalty ($\gamma^{+}$) increases (Figure 6).
As depicted in Figure 6(a), at $\gamma^{+}=$ 2.5 , average objectives are approximately 0.92 for C‑SAA‑CLF, 0.95 for C‑RO‑B‑CLF, and 1.00 for C‑RO‑D‑CLF. As $\gamma^{+}$ rises to 40, these values decline in parallel—C‑SAA‑CLF to 0.63, C‑RO‑B‑CLF to 0.67, and C‑RO‑D‑CLF to 0.73. Throughout this range, C‑SAA‑CLF consistently achieves the lowest average objective, followed by C‑RO‑B‑CLF and then C‑RO‑D‑CLF.
Worst‑case objectives (Figure 6(b)) underscore the robustness of the C‑RO methods: C‑RO‑D‑CLF consistently yields the strongest tail protection, dropping from 0.88 at $\gamma^{+}=$ 2.5 to 0.72 at $\gamma^{+}=$ 40. C‑RO‑B‑CLF falls from 0.885 to 0.775, while C‑SAA‑CLF declines more steeply from 0.925 to 0.73. Notably, at $\gamma^{+}=$ 40, C‑SAA‑CLF nearly matches C‑RO‑D‑CLF, demonstrating that C-SAA can approach robust performance under extreme penalty scenarios.
This robustness manifests in cumulative lateness (Figure 6(c)). Both C‑RO variants tightly control lateness with minimal sensitivity to $\gamma^{+}$, whereas C‑SAA‑CLF’s lateness drops substantially from 0.67 at $\gamma^{+}=$ 2.5 to 0.46 at $\gamma^{+}=$ 40. These robustness gains incur higher expected fulfillment costs (Figure 6(d)). C‑RO‑D‑CLF’s average cost increases from 1.07 to 1.20, and C‑RO‑B‑CLF’s from 1.01 to 1.05, and C‑SAA‑CLF from 0.96 to 1.05. Thus, while all methods incur higher costs under stricter penalties, C‑RO‑B‑CLF exhibits the smallest relative cost increase, C‑SAA‑CLF and C‑RO‑D‑CLF have more aggressive increases in costs due to more conservative decisions.
A final strategic consideration lies in selecting $\gamma^{+}$ to reflect firm priorities. Lower $\gamma^{+}$ emphasizes immediate cost savings—maximizing short‑term profits—whereas higher $\gamma^{+}$ prioritizes on‑time delivery, fostering customer satisfaction, loyalty, and long‑term value. The CSOFP framework encodes this trade‑off directly via $\gamma^{+}$; robust formulations (C‑RO‑B‑CLF and C‑RO‑D‑CLF) further reduce the need for precise $\gamma^{+}$ tuning by delivering stable, worst‑case performance across penalty levels, while C‑SAA‑CLF can leverage high $\gamma^{+}$ most aggressively when strict timeliness is paramount.
(a) Performance in Average Realized Objective Value.
(b) Performance in Worst-Case Realized Objective Value.
(c) Performance in Cumulative Lateness.
(d) Performance in Fulfillment costs.
Figure 6: C-RO vs. C-SAA Across Different Late Delivery Penalties.
@@ 6.7.1 Flexibility of C-RO-B
The C‑RO models effectively hedge against worst‐case delivery deviations and potential misspecifications in the prediction models. However, this robustness incurs more conservative decisions and higher fulfillment costs. The C‑RO‑B variants retain the distribution‐free guarantees while introducing two tuning parameters—prediction‐interval coverage and robust uncertainty budget—that directly control conservativeness. Lower coverage levels and smaller budgets yield less conservative, more cost‐efficient solutions (at the expense of risk protection), whereas higher values emphasize hedging against extreme deviations by sacrificing average performance.
Figures 7 and 8 summarize how these parameters affect solution quality under both C‑RO‑B‑CLF and C‑RO‑B‑QRF. As the prediction interval coverage increases from 70% to 90%. In Figure 7(a), increasing coverage from 70% to 90% reduces the average objective to about 0.84 (C‑RO‑B‑CLF) and 0.86 (C‑RO‑B‑QRF), an approximately 4% improvement relative to 70% coverage. This is consistent with the concurrent decline in cumulative lateness shown in Figure 7(c). However, pushing coverage beyond 95% reverses this trend: at 99% coverage the average objective rises sharply due to the cost of extreme conservatism and, for C‑RO‑B‑CLF, degraded prediction accuracy (reflected in a spike in lateness). By contrast, C‑RO‑B‑QRF continues to reduce lateness at 99%, illustrating its superior tail modeling. In terms of worst‐case performance (Figure 7(b)), both methods improve with coverage, but C‑RO‑B‑CLF plateaus after 95%, whereas C‑RO‑B‑QRF continues to improve at 99%.
The effect of the robust budget, controlled by the scaling factor $\eta$, appears in Figure 8. Increasing $\eta$ from zero to 0.01 leads to a steep drop in the average objective for both methods (Figure 8(a)). Both method reach their minimum values at $\eta=0.01$. At the same time, the worst-case objective follows a similar pattern (Figure 8(b)): it falls sharply up to $\eta=0.1$, then edges back up for both methods as the budget grows. This trend is also seen in cumulative lateness, which drops dramatically until $\eta=0.1$ and then levels off (Figure 8(c)). Together, these trends show that moderate budgets capture most of the gains in both average and worst case performance, while further increases deliver diminishing returns.
(a) Performance in Avg. Obj.
(b) Performance in Worst-Case Obj.
(c) Performance in Cumulative Lateness.
@@
(a) Performance in Avg. Obj.
(b) Performance in Worst-Case Obj.
(c) Performance in Cumulative Lateness.
@@ 6.8 Impact of Consolidation Factor
This section provides a discussion on the impact of item consolidation. CSOFP relies on the coefficients $\beta_{k}$ to control the incentive to batch items. The industrial partner does not know the exact percentage of discount to be offered for item consolidation, which may depend on a variety of convoluted factors. Therefore, this section performs sensitivity analysis on a range of possible discount factors (from 0.1 to 0.7).
Figure 9 compares the performance of three leading methods—C-SAA-CLF, C-RO-B-CLF, and C-RO-D-CLF across average and worst-case realized objective, fulfillment cost, and consolidation level. Consolidation level is defined as the percentage of units batched (sourced and shipped by the same location–carrier pair) in each order.
As shown in Figure 9(a), deeper discounts lead each method to exploit batching incentives more aggressively. C-SAA-CLF consistently attains the lowest average objective across all discount factors, declining from approximately 0.875 at $\beta=$0.1 to about 0.785 at $\beta=$0.7. C‑RO‑B‑CLF follows closely, from 0.895 down to 0.805, while C‑RO‑D‑CLF remains the most conservative, decreasing from 0.955 to 0.905.
The worst‑case objective exhibits a more differentiating pattern (Figure 9(a)). Both C‑RO‑D‑CLF and C‑RO‑B‑CLF are less susceptible to change in the discount, whereas C‑SAA‑CLF starts at 0.842 and falls to 0.802 under high discounts.
@@
As expected, all three methods exhibit lower average objective values due to lower fulfillment costs as the discount factor increases (see Figure 9(c)). This aligns with the intuition that higher discounts encourage more cost-efficient fulfillment strategies. The relative decrease in fulfillment costs across the methods can be attributed to their higher consolidation levels as the discount factor increases, as shown in Figure 9(d). Notably, C-SAA-CLF responds most sharply, with its consolidation rate rising the most rapidly as the discount grows, which is consistent with its pronounced improvement in worst‑case objective performance.
(a) Performance in Average Realized Objective Value.
(b) Performance in Worst-Case Realized Objective Value.
(c) Performance in Fulfillment Costs.
(d) Performance in Consolidation Level.
@@
Figure 9: Performance of CSO Methods Across Different Consolidation Discount Factors.
6.9 Perturbation Analysis under Varying Simulation Environments
To assess the robustness of the proposed CSO methods, this section evaluates performance under varying simulation environments by introducing systematic perturbations to the simulated delivery time deviation distributions. These perturbations represent moderate to adversarial divergence from the test distribution, providing insight into the methods’ stability under distributional shifts.
Two types of perturbations are considered: random uniform noise and delay-specific bias applied to the simulation probability distributions.
Random Uniform Noise Perturbation
Let the original simulation distribution for each test instance be denoted by $P^{sim}=\{p_{i}\}_{i=1}^{C}$. For each class $i\in[C]$, a random noise term $\delta_{i}\sim\mathcal{U}(-\epsilon,\epsilon)$ is added to the probability $p_{i}$, resulting in: $\tilde{p}_{i}=p_{i}+\delta_{i}.$ Negative values are clipped: $\tilde{p}_{i}=\max(\tilde{p}_{i},0)$, and the resulting vector is normalized to form a valid probability distribution: $\tilde{p}_{i}=\tilde{p}_{i}/\sum_{j}\tilde{p}_{j}$.
Figure 10 reports the average and worst‐case realized objectives as the noise level $\epsilon$ increases. In the average case (Figure 10(a)), both C‑RO variants degrade more gradually than C‑SAA, with C‑RO‑B‑CLF becoming the top performer beyond $\epsilon=10$. In the worst‐case scenario (Figure 10(b)), while all methods experience performance degrade, the ranking remains unchanged: C‑RO‑D achieves the lowest realized objective, followed by C‑RO‑B, and then C‑SAA. These results indicate the superior robustness of the C‑RO methods under uniform noise.
Delay-Specific Bias Perturbation
Define the set of late‐delay classes $C_{\mathrm{late}}=\{\,i:\xi_{i}>0\}$. A multiplicative bias $\eta$ is applied to those classes: $\tilde{p}_{i}=p_{i}\cdot(1+\eta).$ The probabilities for classes not in $C_{\mathrm{late}}$ remain unchanged. The resulting vector is clipped to the interval $[0,1]:$ $\tilde{p}_{i}=\max(\min(\tilde{p}_{i},1),0)$, followed by normalization.
Figure 11 shows performance metrics for bias levels up to 250%. In the average case (Figure 11(a)), all CSO methods demonstrate improved performance relative to the Greedy baseline, which deteriorates rapidly. This trend is evidenced by the overall decrease in normalized objective values as the bias parameter $\eta$ increases. Notably, C‑RO‑B overtakes C‑SAA once $\eta$ exceeds 150%, highlighting its robustness under adversarial delay distortions. In the worst‐case objective (Figure 11(b)), C‑RO‑D consistently achieves the best performance, followed by C‑RO‑B‑CLF. Overall, the C‑RO methods exhibit lower variability in performance compared to C‑SAA, indicating enhanced stability under adversarial delay conditions.
(a) Change in Average Objective.
(b) Change in Worst-Case Objective.
Figure 10: Performance Across Different Random Uniform Noise Perturbations.
(a) Change in Average Objective.
(b) Change in Worst-Case Objective.
Figure 11: Performance Across Different Delay-Specific Bias Perturbations.
6.10 Computational Scalability of the CSO Methods
The computational scalability of the proposed framework was evaluated by measuring both the change in computation time (model construction + solve time) and solution quality as instance size grows. Each CSOFP instance is characterized by the number of items $|\mathcal{I}|$ and the number of location–carrier pairs (fulfillment options) $|\mathcal{K}|\times|\mathcal{L}|$. In practice, the eligibility constraint (Constraint (3)) is used to prune infeasible location–carrier pairs in advance. Then, $(\mathcal{K}^{\prime},\mathcal{L}^{\prime})\;=\;\bigl{\{}(k,\ell)\in\mathcal{K}\times\mathcal{L}\,\bigm{|}\,\exists\,i\in\mathcal{I}:\;e_{ik\ell}=1\bigr{\}}$ yields the final set of eligible fulfillment options and thus the effective problem size.
Figure 12(a) shows computation time as the number of eligible fulfillment options increases. Only the MC‑CLF variants of C‑RO and C‑SAA are displayed, since their QRF counterparts exhibit nearly identical runtimes. All three methods achieve optimality on every instance within minutes, confirming their practical applicability. In particular, C‑RO‑B solves most instances in under one second (never exceeding three seconds), and C‑RO‑D solves most in under ten seconds.
These timing differences reflect each method’s formulation. Compared to the nominal problem, the robust counterpart in C‑RO‑B introduces only continuous variables whose count grows linearly with $|\mathcal{K}^{\prime}|$ and $|\mathcal{L}^{\prime}|$. By contrast, C‑RO‑D adds $N$ extra constraints from sampling, and C‑SAA requires solving $Q$ independent problems whose objective function complexity scales with the sample size $N_{1}$. Parallel computation of the $Q$ subproblems can further reduce C‑SAA’s runtime.
Since single-item orders account for the majority of order volume, multi-item orders (what the methods are designed for) can be routed to a dedicated sourcing engine without disrupting standard operations. Under normal load, any of the three methods may be applied; under peak loads—such as holiday surges—C‑RO‑B is recommended to preserve fast solve times while maintaining high solution quality. Concurrent batches of orders can be processed in parallel when computational resources permit.
Figure 12(b) reports average and worst-case objective values for small ($\leq 400$ eligible options), medium ($\leq 1000$ eligible options), and large ($\leq 3000$ eligible options) instances. Average performance remains uniformly high across all size categories, whereas worst-case deviations grow with instance size. Collectively, these results demonstrate near-linear scalability in computation time, consistent high solution quality, and robustness across problem sizes, providing empirical evidence of the framework’s practical scalability.
(a) Change in the Computation Time vs. the Number of Eligible Fulfillment Options.
(b) Objective Value by Instance Categories.
Figure 12: Computational Scalability Analysis.
6.11 Managerial and Operational Implications
The proposed fulfillment framework offers actionable guidance for practitioners seeking to align sourcing strategies with demand variability, service-level requirements, and computational constraints. Depending on the operating context, different methods are preferable. When demand is stable and the goal is to minimize average fulfillment cost, C-SAA-CLF provides the most cost-effective solution. Under moderate runtime constraints, C-RO-B-CLF offers a balanced trade-off between efficiency and robustness. In contrast, during high-stakes periods such as holiday promotions or flash sales—where the cost of service-level violations is high—C-RO-D-CLF and C-RO-B-QRF (with 99% coverage) are better suited to mitigate tail risks and ensure reliability. This scenario-driven approach allows e-commerce platforms to dynamically adapt fulfillment strategies to shifting conditions while leveraging a unified optimization infrastructure. As retailers increasingly face variable demand patterns and rising customer expectations, the ability to flexibly match fulfillment strategy to context becomes a key operational advantage.
Empirical evaluations indicate that the proposed methods can reduce combined fulfillment cost and delivery-timeliness penalties by an average of up to 18% per order under the default settings. To illustrate the potential impact, consider a retailer operating at a scale comparable to the industrial partner, with per-order costs ranging from \$5 to \$30 and monthly volumes of 0.1 to 1 million multi-item orders. Under these scenarios, the corresponding annual savings could range from several million to tens of millions of dollars. Moreover, this framework can be readily extended to single-item orders, which could yield further savings with minimal incremental investment.